# PR2: A Physics- and Photo-realistic Humanoid Testbed with Pilot Study in Competition

### v1

@@ III-C The Virtual Testbed
To ensure stability during simple standing as well as complex tasks such as traversing rough terrains or performing loco-manipulation, three basic controllers are provided for the KUAVO durimh the competition: (i) a standard walking controller that accepts body velocity commands (3-DoFs) for Tasks 1, 2, 4, and 6; (ii) a dynamic jumping controller based on a centroidal dynamics model [29], which takes inputs for jump time, velocity, and height for Task 3; and (iii) a step walking controller based on the Zero Moment Point (ZMP) method [30] for Task 5. The arm controller is not provided and must be developed by the teams to successfully complete Tasks 2, 5, and 6. Critical implementation details for each task are outlined below:
@@
Task 2: The users could turn the valve (diameter around ) either by hitting it with the robot arm’s distal end or by establishing a fixed attachment with the valve when the distal end reaches the enlarged valve volume (see Fig. 3(a)) by sending a “pick” command; this allows the valve revolves together with the robot’s locomotion.
@@
Task 6: Robot would need visual perception (to recognize scenes and objects), language understanding (to translate questions and instructions into actions), and navigation in complex environments (to move and find things). When the user sends a pick command, a fixed joint will be created only if the Euclidean distance between the end-effector and the object to be picked is less than , as shown in this spherical region.
For each task, the robot’s initial pose is randomly sampled within a designated start region. More sophisticated interactions, such as grasping an object through a multi-finger gripper are achievable using finer-grained collision-checking methods.
IV Results
@@
IV-A Platform Capability
@@
Kinodynamic Locomotion: Leveraging the underlying PhysX engine and the Issac Sim APIs, PR2 enables the execution of generated kinodynamic motion plans. With built-in controllers, Fig. 4a shows the robot performing vertical jumping to a moderate height and to an extreme height that is pushing the controller to its limit. Such jumping trajectories are naturally kinodynamic due to the launching, flight, and landing phases during the jump all pose unique challenges in dynamics planning and control, and all require high-fidelity dynamical simulation to capture and reflect such processes.
@@
Specifically, (i) after generating trajectories for the center of mass (CoM) and centroidal angular momentum (CAM) within the hardware capabilities, PR2 needs to accurately compute the robot’s dynamical state evolvement. (ii) In the flight phase, the platform must determine the controller’s trajectory tracking performance to reflect the disturbances from execution errors. (iii) As the robot lands from a higher jump, it experiences larger impact forces. Thus, PR2 must not only reflect different landing impacts but also propagate such impacts to the landing controller to determine robot stability. Fig. 4a demonstrates the execution of two significantly different jump trajectories, wherein the red one takes longer for the controller to stabilize the robot after landing. The capability to reflect the subtle differences in locomotion is a critical aspect of a humanoid testbed.
@@
Physics-based Interaction: Together with the realistic simulation of robot locomotion, PR2 further enables humanoid’s physics-based interaction with the environment during manipulation. Similar to the jumping example, Fig. 4b shows how the reactive force caused by the target valve influences robot motions in loco-manipulation. With a naive controller, the reactive force causes a dramatic torso pitch, whereas a more stable controller could successfully reject the reactive force and maintain body balance. While humanoid robots’ bipedal locomotion
@@
VLA Interface: We introduce customized APIs that enable users to interact with large language models (LLMs) and vision-language models (VLMs), providing a new approach to exploring AI and robotics. In Task 6, users receive visual observations and language instructions to identify goals, formulate plans, and compute robot actions (e.g. joint positions or efforts) at each timestep. The target object in the language instruction is randomly selected from a set in PR2, and the robot is initialized at a random location in the same room as the target object.
@@
Fig. 5 illustrates a typical example where the instruction is put the [pencil on the desk] on the coffee table in the living room. The task involves both object detection (identifying the pencil) and scene understanding (locating the coffee table). Typically, the pencil is in a different room from the coffee table, requiring the user to generate high-level navigation and motion plans to guide the robot in avoiding obstacles and executing the necessary actions. The agent must take the pick action when it believes it is close enough to the target object. If the distance is within a specified threshold, a fixed joint is created to attach the target object to the robot’s end effector. Upon reaching the coffee table, the agent must take the release action to complete the task.
IV-B Team Performance Analysis
@@
Fig. 6 presents the score distribution percentages across the six tasks. Each bar representing a task is divided into colored segments corresponding to different score ranges.
@@
Fig. 7 further presents the individual time consumption for each score across all tasks, with each score represented by a distinct box plot that shows the distribution of times achieved by participants. The overall plot reveals varying levels of difficulty and completion times for each task. Generally, the robot requires more time as it becomes more involved in the task to achieve higher scores. Task 6 exhibits the widest range of scores and the most balanced distribution, indicating diverse levels of student performance. This task saw some students excelling, while others struggled, as evidenced by the highest proportion of participants scoring 0 and the second highest scoring 25. The presence of outliers across all tasks suggests that while most students completed within a typical range, some took significantly longer.
By categorizing the detailed score rubric into three categories—locomotion (60%), manipulation (30%), and perception (10%)—as shown in Tab. II, we analyzed the average points participants scored (see Fig. 8). The results indicate that participants performed relatively well in locomotion, earning nearly half of the available points (27.55 out of 60), reflecting the emphasis this area has received in humanoid literature. However, they encountered significantly greater challenges in tasks involving manipulation and perception, scoring less than 30% and 20% of the available points in these categories, respectively. Fig. 9 shows some typical failure cases in manipulating the valve (Task 2) and perception (Task 6).
IV-C Lesson Learned
@@
Threeway tradeoff among rendering quality, simulation precision, and computing resource.
@@
Moreover, only a small percentage of participants had access to high-end consumer-grade GPUs like the RTX 40 Series, which can cost up to . To accommodate these limitations, we implemented new APIs that allowed participants to lower rendering quality and increase the simulation time-step. However, these adjustments inevitably reduced the precision and accuracy of the simulation. Balancing rendering quality, simulation precision, and computing resources is a crucial design consideration for promoting more accessible and equitable education opportunities in humanoid robotics.
@@
Model-based controller vs. Model-free policy. Although the competition provided basic controllers for robot stabilization and tracking, some participants opted to develop and train RL-based policies instead. While both approaches can effectively generate robot motions and are encouraged, they introduce significant challenges in (i) creating a robust testbed that supports both model-based and model-free methods, potentially combining elements of them in hybrid approaches, and (ii) ensuring fair evaluation as RL often require significant computational resources for training.
@@
The necessity and challenges in integrating AI-related topics. Participants generally performed well in locomotion-related tasks with the provided controllers, as shown in Fig. 8. However, significant performance drops were observed when tasks incorporated elements of manipulation and visual or language perception, as reflected by lower average scores. Specifically, difficulties were most pronounced in Task 3 and 6, which required high-level planning and the integration of AI techniques. These results underscore the need for future research and education in humanoid robotics to encompass a broader range of skills and topics, addressing the challenges that arise when integrating complex AI components with robot tasks.
V Conclusion and Looking Forward
@@
In this paper, we presented the development of PR2, a testbed that combines physics-realistic simulation with photo-realistic scene rendering. By integrating high-quality scene rendering with high-fidelity simulation of robot dynamics, PR2 offers a new platform for exploring full-size humanoid robot locomotion, manipulation, and perception, addressing critical aspects for future humanoid applications. The deployment of PR2 in an online college student competition attracted hundreds of participants, revealing the challenges junior students faced in combining manipulation and perception in humanoid robot locomotion and emphasizing the need for future research and education to focus on these complex integrations. We have released an upgraded version of the PR2 testbed, which we believe will significantly benefit future research and education in humanoid robotics, particularly for those who previously lacked access. Moving forward, our priorities include incorporating a built-in RL training scheme alongside planning and control methods, enhancing scene diversity through scene synthesis techniques [31], and refining fine-grained grasping capabilities for loco-manipulation tasks in future PR2 releases.
@@
Acknowledgement: This work was supported in part by the National Natural Science Foundation of China (No. 62403064, 62403063). We thank the engineering team from Leju Robotics for technical support, Ms. Zhen Chen from BIGAI for refining the figures, and all colleagues from the BIGAI TongVerse project for fruitful discussions and help on simulation developments.

### v2

@@ 3.3 The Virtual Testbed
To ensure stability during simple standing as well as complex tasks such as traversing rough terrains or performing loco-manipulation, three basic controllers are provided for the KUAVO during the competition: (i) a standard walking controller that accepts body velocity commands (3-DoFs) for Tasks 1, 2, 4, and 6; (ii) a dynamic jumping controller based on a centroidal dynamics model he2024cdm, which takes inputs for jump time, velocity, and height for Task 3; and (iii) a step walking controller based on the Zero Moment Point (ZMP) method scianca2020mpc for Task 5. The arm controller is not provided and must be developed by the teams to complete Tasks 2, 5, and 6 successfully. Critical implementation details for each task are outlined below:
@@
Task 2: The users could turn the valve (diameter around ) either by hitting it with the robot arm’s distal end or by establishing a fixed attachment with the valve when the distal end reaches the enlarged valve volume (see Figure 5(b)) by sending a “pick” command; this allows the valve revolves together with the robot’s locomotion.
@@
Task 6: The robot would need visual perception (to recognize scenes and objects), language understanding (to translate questions and instructions into actions), and navigation in complex environments (to move and find things). When the user sends a pick command, a fixed joint will be created only if the Euclidean distance between the end-effector and the object to be picked is less than , expressed as a spherical region as shown in Figure 5(a).
For each task, the robot’s initial pose is randomly sampled within a designated start region. Although the objective interactions in task 2 and task 6 have been simplified for the college competition, the underlying PhysX engine is fully capable of supporting more advanced manipulation tasks. For example, complex interactions, such as grasping objects with a multi-finger gripper, can be achieved by utilizing finer-grained collision-checking and more sophisticated control strategies. It is important to note that while these capabilities exist within our simulation framework, the explicit demonstration remains one of our ongoing efforts in future PR2 releases.
4 Results
@@
4.1 Platform Capability
4.1.1 Kinodynamic Locomotion
@@
Leveraging the underlying PhysX engine and the Isaac Sim APIs, PR2 enables the execution of generated kinodynamic motion plans. With built-in controllers, Figure 6(a) shows the robot performing vertical jumping to a moderate height and to an extreme height that is pushing the controller to its limit. Such jumping trajectories are naturally kinodynamic due to the launching, flight, and landing phases during the jump all pose unique challenges in dynamics planning and control, and all require high-fidelity dynamical simulation to capture and reflect such processes.
@@
Specifically, (i) after generating trajectories for the center of mass (CoM) and centroidal angular momentum (CAM) within the hardware capabilities, PR2 needs to accurately compute the robot’s dynamical state evolvement. (ii) In the flight phase, the platform must determine the controller’s trajectory tracking performance to reflect the disturbances from execution errors. (iii) As the robot lands from a higher jump, it experiences larger impact forces. Thus, PR2 must not only reflect different landing impacts but also propagate such impacts to the landing controller to determine robot stability. Figure 6(a) demonstrates the execution of two significantly different jump trajectories, wherein the red one takes longer for the controller to stabilize the robot after landing. The capability to reflect the subtle differences in locomotion is a critical aspect of a humanoid testbed.
4.1.2 Physics-based Interaction
@@
Together with the realistic simulation of robot locomotion, PR2 further enables humanoid’s physics-based interaction with the environment during manipulation. Similar to the jumping example, Figure 6(b) shows how the reactive force caused by the target valve influences robot motions in loco-manipulation. With a naive controller, the reactive force causes a dramatic torso pitch, whereas a more stable controller could successfully reject the reactive force and maintain body balance.
4.1.3 VLA Interface
@@
We introduce customized APIs that enable users to interact with large language models (LLMs) and vision-language models (VLMs), providing a new approach to exploring AI and robotics. In Task 6, users receive visual observations and language instructions to identify goals, formulate plans, and compute robot actions (e.g. joint positions or efforts) at each timestep. The target object in the language instruction is randomly selected from a set in PR2, and the robot is initialized at a random location in the same room as the target object.
@@
Figure 7 illustrates a typical example where the instruction is put the [pencil on the desk] on the coffee table in the living room. The task involves both object detection (identifying the pencil) and scene understanding (locating the coffee table). Typically, the pencil is in a different room from the coffee table, requiring the user to generate high-level navigation and motion plans to guide the robot in avoiding obstacles and executing the necessary actions. The agent must take the pick action when it believes it is close enough to the target object. If the distance is within a specified threshold, a fixed joint is created to attach the target object to the robot’s end effector. Upon reaching the coffee table, the agent must take the release action to complete the task.
4.2 Team Performance Analysis
@@
Figure 8 presents the score distribution percentages across the six tasks. Each bar representing a task is divided into colored segments corresponding to different score ranges.
@@
Figure 9 further presents the individual time consumption for each score across all tasks, with each score represented by a distinct box plot that shows the distribution of times achieved by participants. The overall plot reveals varying levels of difficulty and completion times for each task. Generally, the robot requires more time as it becomes more involved in the task to achieve higher scores. Task 6 exhibits the widest range of scores and the most balanced distribution, indicating diverse levels of student performance. This task saw some students excelling, while others struggled, as evidenced by the highest proportion of participants scoring 0 and the second highest scoring 25. The presence of outliers across all tasks suggests that while most students completed within a typical range, some took significantly longer.
By categorizing the detailed score rubric into three categories—locomotion (60%), manipulation (30%), and perception (10%)—as shown in Table 2, we analyzed the average points participants scored (see Figure 10). The results indicate that participants performed relatively well in locomotion, earning nearly half of the available points (27.55 out of 60), reflecting the emphasis this area has received in humanoid literature. However, they encountered significantly greater challenges in tasks involving manipulation and perception, scoring less than 30% and 20% of the available points in these categories, respectively. Figure 11 shows some typical failure cases in manipulating the valve (Task 2) and perception (Task 6).
4.3 Lesson Learned
@@
4.3.1 Trilemma of rendering quality, simulation precision, and computing resource
@@
In the competition, a limited number of participants had access to high-end consumer-grade GPUs, such as the RTX 40 Series, which can cost up to . To address these hardware constraints, we introduced new APIs that enable participants to adjust rendering quality and modify the simulation time-step. Reducing rendering quality can increase the frame rate, ensuring that PR2 can photo-realistically reflect the simulated physics. We conducted benchmarks on a computer equipped with a 13th Gen Intel® Core™ i9-13900KF CPU and an Nvidia 4090 GPU. With a fixed physical simulation time-step of 0.001, Table 4 indicated that the frame rate remains stable until the rendering workload approaches the system’s computational limits. Beyond a resolution of pixels, the frame rate decreases as rendering demands exceed available resources, necessitating adjustments to maintain consistent physics time-steps. Alternatively, increasing the simulation time-step can alleviate computational burdens. Larger time-steps give the physics engine coarser time slices to compute interactions, including the integration of forces and motion, leading to numerical errors and degraded fidelity. Balancing rendering quality, simulation precision, and computing resources is a crucial design consideration for promoting more accessible and equitable education opportunities in humanoid robotics.
| Resolution (pixels) | 640480 | 1024768 | 1280720 | 16001200 |
| Average FPS | 17.363 | 17.683 | 17.361 | 16.263 |
| Resolution (pixels) | 19201080 | 25601440 | 34401440 | 38402160 |
| Average FPS | 14.918 | 15.033 | 13.973 | 13.697 |
4.3.2 Model-based controller v.s. Model-free policy
@@
Although the competition provided basic controllers for robot stabilization and tracking, some participants opted to develop and train RL-based policies instead. While both approaches can effectively generate robot motions and are encouraged, they introduce significant challenges in (i) creating a robust testbed that supports both model-based and model-free methods, potentially combining elements of them in hybrid approaches, and (ii) ensuring fair evaluation as RL often require significant computational resources for training.
4.3.3 The necessity and challenges in integrating AI-related topics
@@
Participants generally performed well in locomotion-related tasks with the provided controllers, as shown in Figure 10. However, significant performance drops were observed when tasks incorporated elements of manipulation and visual or language perception, as reflected by lower average scores. Specifically, difficulties were most pronounced in Task 3 and 6, which required high-level planning and the integration of AI techniques. These results underscore the need for future research and education in humanoid robotics to encompass a broader range of skills and topics, addressing the challenges that arise when integrating complex AI components with robot tasks.
4.3.4 Competition with Physical Robots
The top 10 teams from the simulation competition were invited to participate in the physical humanoid robot competition held in August 2024. Five trials were designed based on the simulated tasks in PR2, with modifications to accommodate physical robots (see Figure 12). For the valve-turning and button-pressing tasks, the robot needed to locate the target object, approach it, and use its arm to either rotate a horizontal valve 180 degrees or press a button. In the uneven terrain walking task, the robot had to maintain stability while navigating rough surfaces. The door-opening task required the robot to grasp and rotate a doorknob. In the indoor object search task, an audio cue indicated the target object the robot needed to fetch, incorporating challenges in speech recognition, object detection, locomotion, and grasping. Each team had one hour for competition and participants could complete the tasks in any order, but bonus points were awarded for completing tasks consecutively.
Several sim-to-real gaps were identified during the competition. Accurately replicating the damping and friction characteristics of both the robot’s actuators and the valves used in the physical competition within the simulation environment proved challenging. This discrepancy can lead to differences in robot performance between simulated and real-world scenarios. Additionally, the simulation environment often fails to capture the complex lighting variations present in real-world settings, affecting the robot’s vision system and its ability to accurately perceive and interact with objects. Despite the substantial domain gap between simulation and real-world execution, PR2 remains an effective testbed for participants to become familiar with the robotic platform and build a foundation for the competition. On average, participants scored , , , and in the valve-turning, button-pressing, uneven terrain walking, and door-opening tasks, respectively—significantly higher than the average performance in simulation competition. However, the indoor object search task saw an average score of only , highlighting the challenges of integrating locomotion with manipulation and visual/audio perception.
5 Conclusion and Looking Forward
@@
In this paper, we presented the development of PR2, a testbed that combines physics-realistic simulation with photo-realistic scene rendering. By integrating high-quality scene rendering with high-fidelity simulation of robot dynamics, PR2 offers a new platform for exploring full-size humanoid robot locomotion, manipulation, and perception, addressing critical aspects for future humanoid applications. The deployment of PR2 in an online collegiate competition attracted hundreds of participants, revealing the challenges junior students faced in combining manipulation and perception in humanoid robot locomotion and emphasizing the need for future research and education to focus on these complex integrations. We have released an upgraded version of the PR2 testbed, which we believe will significantly benefit future research and education in humanoid robotics, particularly for those who previously lacked access. Moving forward, our priorities include incorporating a built-in RL training scheme alongside planning and control methods, enhancing scene diversity through scene synthesis techniques yang2024physcene, and refining fine-grained grasping capabilities for loco-manipulation tasks in future PR2 releases.
Acknowledgement:
@@
This work was supported in part by the National Natural Science Foundation of China (No. 62403064, 62403063). We thank the engineering team from Leju Robot for technical support, Ms. Zhen Chen from BIGAI for refining the figures, and all colleagues from the BIGAI TongVerse project for fruitful discussions and help with simulation developments.